# Discrete State Diffusion Models: A Sample Complexity Perspective

### v1

@@ Abstract
**footnotetext: Equal contribution
The authors are with Purdue University, West Lafayette IN 47907, USA.
@@ 1 Introduction
Diffusion Models (Sohl-Dickstein et al. (2015); Song et al. (2021); Ho and Salimans (2022)) have gained significant attention lately due to their empirical success in various generative modeling tasks across computer vision and audio generation (Ulhaq and Akhtar (2022); Bansal et al. (2023)), text generation (Li et al. (2022)), sequential data modeling (Tashiro et al. (2021)), policy learning (Chen et al. (2024)), life sciences (Jing et al. (2022); Malusare and Aggarwal (2024)), and biomedical image reconstruction (Chung and Ye (2022)). They model an unknown and unstructured distribution through forward and reverse stochastic processes. In the forward process, samples from the dataset are gradually corrupted to obtain a stationary distribution. In the reverse process, a well-defined noisy distribution is used as the initialization to iteratively produce samples that resemble the learnt distribution using the score function. Most of the earlier works have been on continuous-state diffusion models, where the data is defined on the Euclidean space $\mathcal{R}^{d}$, unlike the discrete data, which is defined on $\mathcal{S}^{d}$, where $\mathcal{S}$ is the discrete set of values each component can take and $d$ is the number of components or dimension of the data. Given the growing number of applications of generative AI with discrete data, such as text generation (Zhou et al. (2023)) and summarization (Dat et al. (2025)), combinatorial optimization (Li et al. (2024)), molecule generation and drug discovery (Malusare and Aggarwal (2024)), protein and DNA sequence design (Alamdari et al. (2023); Avdeyev et al. (2023)), and graph generation (Qin et al. (2024)), there has been an increased interest in diffusion with discrete data. For a more complete review of related literature, one may refer to Ren et al. (2025b).
@@
  • First sample complexity bounds for discrete-state diffusion models To the best of our knowledge, this is the first work to present rigorous sample complexity analyses under reasonable and practical assumptions for discrete-state diffusion models. We derive an order optimal bound of $\widetilde{\mathcal{O}}(\epsilon^{-2})$, and thus providing deeper insight into the sample efficiency of discrete-state diffusion.
  • Principled error decomposition We propose a structured decomposition of the score estimation error into approximation, statistical, optimization, and clipping components, enabling the characterization of how each factor contributes to sample complexity.
@@ 2.1 Discrete-state Diffusion Models
Discrete-state diffusion models were initially studied as discrete-time Markov chains, which train and sample the model at discrete time points. Austin et al. (2021) first proposed Discrete Denoising Diffusion Probabilistic Models (D3PMs) as the discrete-state analog of the “mean predicting” DDPMs (Ho et al. (2020)). These models were largely applied to fields other than language due to empirical challenges (Lou et al. (2024)). Although these models offer various practical sampling algorithms, they lack flexibility and theoretical guarantees (Zhang et al. (2025)). Campbell et al. (2022) first introduced the continuous-time (CTMC) framework, similar to the NCSNv3 by Song et al. (2021). The time-reversal of a forward CTMC is fully determined by its rate matrix and the forward marginal probability ratios, which together define the discrete-state score function.
@@ 2.2 Convergence and Theoretical Analysis of Discrete-state Diffusion Models
While convergence analysis for continuous-state diffusion models is well-established, theoretical understanding in the discrete setting remains comparatively underdeveloped. Campbell et al. (2022) provided one of the earliest error bounds for discrete-state diffusion by analyzing a $\tau$-leaping sampling algorithm under total variation distance, assuming bounded forward probability ratios and $L^{\infty}$ control on the rate matrix approximation. However, the bound exhibits at least quadratic growth in the dimension $d$. In contrast, Chen and Ying (2024) introduced a score-based sampling method for discrete-state diffusion models by simulating the reverse process via CTMC uniformization on the binary hypercube, where forward transitions involve independent bit flips. Their method yields convergence guarantees and algorithmic complexity bounds that scale nearly linearly with $d$, comparable to the best-known results in the continuous setting (Benton et al. (2024)). Zhang et al. (2025) and Ren et al. (2025a) perform KL divergence error analysis by decomposing into truncation, discretization, and approximation error, and provide iteration complexities of sampling algorithms.
@@ 3.1 Discrete-state diffusion processes
The probability distributions are considered over a finite support $\mathcal{X}=[S]^{d}$, where each sequence $x\in\mathcal{X}$ has $d$ components drawn from a discrete set of size $S$. The forward marginal distribution $p_{t}$ can be represented as a probability mass vector $p_{t}\in\Delta^{N}$, where $\Delta^{N}$ is the probability simplex in $\mathbb{R}^{N}$, and $N=S^{d}$ is the cardinality of the discrete-state space. The forward dynamics of this distribution are governed by a continuous-time Markov process described by the Kolmogorov forward equation (Campbell et al. (2022)).
$\frac{dp_{t}}{dt}=Q_{t}^{\top}p_{t},\;p_{0}\approx p_{\text{data}}\;$ (1)
Here, $Q_{t}\in\mathbb{R}^{N\times N}$ is a time-dependent generator (or rate) matrix with non-negative off-diagonal entries and rows summing to zero. This ensures conservation of total probability mass as the distribution evolves over time. We assume a finite uniform departure rate as required by the uniformization algorithm, which bounds the state-dependent exit rates by a single global rate (Zhang et al. (2025); Chen and Ying (2024)).
@@
To approximate this continuous evolution in practice, a small-step Euler discretization is applied. The resulting transition probability for a step of size $\Delta t$ is given by (Zhang et al. (2025); Lou et al. (2024)).
$p(x_{t+\Delta t}=y\mid x_{t}=x)=\delta{\{x,y\}}+Q_{t}(x,y)\Delta t+o(\Delta t)$ (3)
Equation (3) captures the probability of transitioning from state $x$ to $y$, $Q_{t}(x,y)$ governing the instantaneous transition rate and $p(x_{t+\Delta t}=y\mid x_{t}=x)$ is the infinitesimal transition probability of being in state $y$ at time $t+\Delta t$ given state $x$ at time $t$. The process also admits a well-defined reverse-time evolution (Kelly (2011)), with $p_{T-t}$ the time-reversed marginal and is characterized by another diffusion matrix, $\overline{Q}_{t}$, as follows:
$\frac{dp_{T-t}}{dt}$ $=\overline{Q}_{T-t}^{T}p_{T-t},\;\overline{Q}_{t}(x,y)=\tfrac{p_{t}(y)}{p_{t}(x)}Q_{t}(y,x)\;$ (4)
@@
$\mathcal{L}_{\text{SE}}(\hat{s}_{\theta})$ $=\int_{0}^{T}\mathbb{E}_{x_{t}\sim q_{t}}\sum_{y\neq x_{t}}Q_{t}(y,x_{t})\,D_{I}\!\left(s_{t}(x_{t})_{y}\,\|\,\hat{s}_{\theta,t}(x_{t})_{y}\right)\,dt$ (5)
which is characterized by the Bregman divergence $D_{I}(\cdot)$ unlike the $L^{2}$ loss widely adopted in the continuous case. The Bregman divergence $D_{I}(\cdot)$, with respect to the negative entropy function $I(\cdot)$, also known as the generalized I-divergence, is defined as:
@@
In Equation (6), $I(\cdot)$ refers to the negative entropy function, where $I(x)=\Sigma_{i=1}^{d}x^{i}logx^{i}$ for a $d$ dimensional data. It is important to note that the Bregman divergence does not satisfy the triangle inequality. We defer more details on the Bregman divergence and convexity of $I(\cdot)$ to Lemma 4 in Appendix B.
@@ 3.2 Forward process
Since directly working with a rate matrix $Q_{t}$ of size $S^{d}\times S^{d}$ is infeasible in high dimension, the model assumes a factorized structure (Zhang et al. (2025); Chen and Ying (2024)). In particular, the conditional distribution of the process can be written as $q_{t|s}(x_{t}\mid x_{s})=\prod_{i=1}^{d}q^{i}_{t|s}(x_{t}^{i}\mid x_{s}^{i})$, so that each coordinate evolves independently as its own CTMC, with forward rate $Q^{\text{tok}}_{t}$. Under this construction, the global rate matrix $Q_{t}$ is sparse: transitions between states are only possible when their Hamming distance is one as adopted by Campbell et al. (2022). A common choice for the per-coordinate dynamics is the time-homogeneous uniform flip rate, given by
@@
Combining these coordinate-wise generators produces the global time-homogeneous rate matrix $Q$, whose entries are given by (Zhang et al. (2025))
@@
The chain evolves by allowing each coordinate to flip independently, but at any given instant only one coordinate flip occurs. Thus transitions always correspond to Hamming distance one. As $t\to\infty$, the marginal $q_{t}$ converges to the uniform distribution $\pi^{d}$ over the entire state space $S]^{d}$.
@@ 3.3 Reverse process
Here, $Q_{T-t}^{\text{tok}}$ is the per-coordinate rate matrix of the forward chain, the indicator $\delta\{x^{\backslash i},\tilde{x}^{\backslash i}\}$ enforces that $x$ and $\tilde{x}$ differ only in coordinate $i$. This construction shows that the reverse process also evolves by flipping one coordinate at a time, but the rates are adjusted by the relative likelihood of the target state compared to the current state under the reverse marginal $q_{T-t}$. This ratio defines the discrete-state score function, which is central to training the reverse-time model. From the structure of the rate matrix as defined above, Equation (5) can be compactly written as follows:
@@ 3.4 Problem Formulation
Let the discrete-state score function be approximated by a parameterized family of functions $\mathcal{F}_{\Theta}=\{\hat{s}_{\theta}:\theta\in\Theta\}$, where each $\hat{s}_{\theta}:[S]^{d}\times[0,T]\to\mathbb{R}^{d(S-1)}$ is typically represented by a neural network with depth $D$ and width $W$. Given i.i.d. samples $\{x_{i}\}_{i=1}^{n}\sim p_{\text{data}}$ from the data distribution, the score network is trained by minimizing the following time-indexed loss:
@@ 4 Methodology
$\mathcal{L}_{k}(\theta)$ $=\int_{kh+\delta}^{(k+1)h+\delta}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\left[D_{I}\left(s_{k}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,k}(\mathbf{x}_{t})\right)\right]dt$ (14)
@@
$\widehat{\mathcal{L}}_{k}(\theta)=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}D_{I}\left(s_{k}(\mathbf{x}_{k,t,i})\,\|\,\hat{s}_{\theta,k}(\mathbf{x}_{k,t,i})\right),\quad t\sim\text{Unif}[kh,(k+1)h]$
@@
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks (Liu et al. (2022)). Prior works in continuous-state diffusion models such as Gupta et al. (2024) and Block et al. (2020) implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in Gupta et al. (2024) and the definition of $\hat{f}$ in Theorem 13 of Block et al. (2020)).This assumption, however, introduces a major limitation for practical implementations, where the exact ERM is not attainable. In contrast, the PL condition allows us to derive sample complexity bounds under realistic optimization dynamics, without requiring exact ERM solutions.
@@
The bounded gradient variance assumption is a standard assumption used in SGD convergence results such as Koloskova et al. (2022) and Ajalloeian and Stich (2020). This assumption holds for a broad class of neural networks under mild conditions, such as Lipschitz-continuous activations (e.g., GELU), bounded inputs, and standard initialization schemes (Allen-Zhu et al. (2019); Du et al. (2019)). The smoothness condition ensures that the gradients of the loss function do not change abruptly, which is crucial for stable updates during optimization. This stability facilitates controlled optimization dynamics when using first-order methods like SGD or Adam, ensuring that the learning process proceeds in a well-behaved manner.
@@
The data distribution $p_{\text{data}}$ has full support on $\mathcal{X}$, and there exists a uniform $B>0$, such that for all $\mathbf{x}\in\mathcal{X}$, $i\in[d]$, and $x^{i}\neq\hat{x}^{i}\in[S]$,

### v2

@@ 1 Introduction
Diffusion Models (Song et al., 2021) have gained significant attention lately due to their empirical success in various generative modeling tasks across computer vision and audio generation (Ulhaq and Akhtar, 2022; Bansal et al., 2023), text generation (Li et al., 2022), sequential data modeling (Tashiro et al., 2021), policy learning (Chen et al., 2024), life sciences (Jing et al., 2022; Malusare and Aggarwal, 2025; Malusare et al., 2025), and biomedical image reconstruction (Chung and Ye, 2022). They model an unknown and unstructured distribution through forward and reverse stochastic processes. In the forward process, samples from the dataset are gradually corrupted to obtain a stationary distribution. In the reverse process, a well-defined noisy distribution is used as the initialization to iteratively produce samples that resemble the learnt distribution using the score function. Most of the earlier works have been on continuous-state diffusion models, where the data is defined on the Euclidean space $\mathcal{R}^{d}$, unlike the discrete data, which is defined on $\mathcal{S}^{d}$, where $\mathcal{S}$ is the discrete set of values each component can take and $d$ is the number of components or dimension of the data. Given the growing number of applications of generative AI with discrete data, such as text generation (Zhou et al., 2024) and summarization (Do et al., 2025), combinatorial optimization (Li et al., 2024), molecule generation and drug discovery (Malusare and Aggarwal, 2025), protein and DNA sequence design Alamdari et al. (2023); Avdeyev et al. (2023), and graph generation Qin et al. (2023), there has been an increased interest in diffusion with discrete data. For a more complete review of related literature, one may refer to Ren et al. (2025b).
@@
  • First sample complexity bounds for discrete-state diffusion models To the best of our knowledge, this is the first work to present rigorous sample complexity analyses under reasonable and practical assumptions for discrete-state diffusion models. We derive an order optimal bound of $\widetilde{\mathcal{O}}(\epsilon^{-2})$, and thus providing deeper insight into the sample efficiency of discrete-state diffusion. In particular we prove that $\widetilde{\mathcal{O}}(\epsilon^{-2})$ samples are enough for discrete-state diffusion models to generate samples from a distribution that is $\epsilon$-close to the target distribution in KL divergence.
  • Principled error decomposition We propose a structured decomposition of the score estimation error into approximation, statistical, optimization, and clipping components, enabling the characterization of how each factor contributes to the overall sample complexity.
@@ 2.1 Discrete-state Diffusion Models
Discrete-state diffusion models were initially studied as discrete-time Markov chains, which train and sample the model at discrete time points. Austin et al. (2021) first proposed Discrete Denoising Diffusion Probabilistic Models (D3PMs) as the discrete-state analog of the “mean predicting” DDPMs (Ho et al., 2020). These models were largely applied to fields other than language due to empirical challenges (Lou et al., 2024). Although these models offer various practical sampling algorithms, they lack flexibility and theoretical guarantees (Zhang et al., 2025). Campbell et al. (2022) first introduced the continuous-time (CTMC) framework, similar to the NCSNv3 by Song et al. (2021). The time-reversal of a forward CTMC is fully determined by its rate matrix and the forward marginal probability ratios, which together define the discrete-state score function.
@@ 2.2 Convergence and Theoretical Analysis of Discrete-state Diffusion Models
While convergence analysis for continuous diffusion models is well-established, theoretical understanding in the discrete setting remains comparatively underdeveloped. Campbell et al. (2022) provided one of the earliest error bounds for discrete-state diffusion by analyzing a $\tau$-leaping sampling algorithm under total variation distance, assuming bounded forward probability ratios and $L^{\infty}$ control on the rate matrix approximation. However, the bound exhibits at least quadratic growth in the dimension $d$. In contrast, Chen and Ying (2024) introduced a score-based sampling method for discrete-state diffusion models by simulating the reverse process via CTMC uniformization on the binary hypercube, where forward transitions involve independent bit flips. Their method yields convergence guarantees and algorithmic complexity bounds that scale nearly linearly with $d$, comparable to the best-known results in the continuous setting (Benton et al., 2024a). Zhang et al. (2025) and Ren et al. (2025a) perform KL divergence error analysis by decomposing into truncation, discretization, and approximation error, and provide iteration complexities of sampling algorithms. However, these analyses assume the score estimation error as simply bounded by a constant ( denoted by $\epsilon_{\mathrm{score}}$), effectively bypassing the score-learning problem.
Several recent works have analyzed the sample complexity of score-based continuous diffusion models. For example, Gupta et al. (2024) presents a bound of $\widetilde{\mathcal{O}}(\epsilon^{-5})$ using a quantile-based reformulation and assuming access to the ERM. In addition, Gaur et al. (2025b) derives a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$ by relaxing the assumption of access to ERM. Extending these results to discrete-state diffusion models, we establish for the first time a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-2})$.
@@ 3.1 Discrete-state diffusion processes
The probability distributions are considered over a finite support $\mathcal{X}=[S]^{d}$, where each sequence $x\in\mathcal{X}$ has $d$ components drawn from a discrete set of size $S$. The forward marginal distribution $q_{t}$ can be represented as a probability mass vector $q_{t}\in\Delta^{N}$, where $\Delta^{N}$ is the probability simplex in $\mathbb{R}^{N}$, and $N=S^{d}$ is the cardinality of the discrete-state space. The forward dynamics of this distribution are governed by a continuous-time Markov process described by the Kolmogorov forward equation (Campbell et al., 2022).
$\frac{dq_{t}}{dt}=Q_{t}^{\top}q_{t},\;q_{0}\approx q_{\text{data}}\;$ (1)
Here, $Q_{t}\in\mathbb{R}^{N\times N}$ is a time-dependent generator (or rate) matrix with non-negative off-diagonal entries and rows summing to zero. This ensures conservation of total probability mass as the distribution evolves over time. We assume a finite uniform departure rate as required by the uniformization algorithm, which bounds the state-dependent exit rates by a single global rate (Zhang et al., 2025; Chen and Ying, 2024).
@@
To approximate this continuous evolution in practice, a small-step Euler discretization is applied. The resulting transition probability for a step of size $\Delta t$ is given by (Zhang et al., 2025; Lou et al., 2024).
$q(x_{t+\Delta t}=y\mid x_{t}=x)=\delta{\{x,y\}}+Q_{t}(x,y)\Delta t+o(\Delta t)$ (3)
Equation (3) captures the probability of transitioning from state $x$ to $y$, $Q_{t}(x,y)$ governing the instantaneous transition rate and $q(x_{t+\Delta t}=y\mid x_{t}=x)$ is the infinitesimal transition probability of being in state $y$ at time $t+\Delta t$ given state $x$ at time $t$. The process also admits a well-defined reverse-time evolution (Kelly, 2011), with $q_{T-t}$ the time-reversed marginal and is characterized by another diffusion matrix, $\overline{Q}_{t}$, as follows:
$\frac{dq_{T-t}}{dt}$ $=\overline{Q}_{T-t}^{T}q_{T-t},\;\overline{Q}_{t}(x,y)=\tfrac{q_{t}(y)}{q_{t}(x)}Q_{t}(y,x)\;$ (4)
@@
$\mathcal{L}_{\text{SE}}(\hat{s}_{\theta})$ (5)
$=\int_{0}^{T}\mathbb{E}_{x_{t}}\sum_{y\neq x_{t}}Q_{t}(y,x_{t})\,D_{I}\!\left(s_{t}(x_{t})_{y}\,\|\,\hat{s}_{\theta,t}(x_{t})_{y}\right)\,dt$
where the expectation is taken over $x_{t}\sim q_{t}$} and is characterized by the Bregman divergence $D_{I}(\cdot)$ unlike the $L^{2}$ loss widely adopted in the continuous case. The Bregman divergence $D_{I}(\cdot)$, with respect to the negative entropy function $I(\cdot)$, also known as the generalized I-divergence, is defined as:
@@
In Equation (6), $I(\cdot)$ refers to the negative entropy function, where $I(x)=\Sigma_{i=1}^{d}x^{i}\log x^{i}$ for a $d$ dimensional data. It is important to note that the Bregman divergence does not satisfy the triangle inequality. We defer more details on the Bregman divergence and convexity of $I(\cdot)$ to Lemma 4 in Appendix B.
@@ 3.2 Forward process
Since directly working with a rate matrix $Q_{t}$ of size $S^{d}\times S^{d}$ is infeasible in high dimension, the model assumes a factorized structure (Zhang et al., 2025; Chen and Ying, 2024). In particular, the conditional distribution of the process can be written as $q_{t|s}(x_{t}\mid x_{s})=\prod_{i=1}^{d}q^{i}_{t|s}(x_{t}^{i}\mid x_{s}^{i})$, so that each coordinate evolves independently as its own CTMC, with forward rate $Q^{\text{tok}}_{t}$. Under this construction, the global rate matrix $Q_{t}$ is sparse: transitions between states are only possible when their Hamming distance is one as adopted by Campbell et al. (2022). A common choice for the per-coordinate dynamics is the time-homogeneous uniform flip rate, given by
@@
Combining these coordinate-wise generators produces the global time-homogeneous rate matrix $Q$, whose entries are given by (Zhang et al., 2025)
@@
The chain evolves by allowing each coordinate to flip independently, but at any given instant only one coordinate flip occurs. Thus transitions always correspond to Hamming distance one. We defer the details on the factorisation and effective score space to Appendix F. As $t\to\infty$, the marginal $q_{t}$ converges to the uniform distribution $\pi^{d}$ over the entire state space $S]^{d}$ (Chen and Ying, 2024; Zhang et al., 2025).
@@ 3.3 Reverse process
Here, $Q^{\text{tok}}$ is the time-homogenous per-coordinate rate matrix of the forward chain, the indicator $\delta\{x^{\backslash i},\tilde{x}^{\backslash i}\}$ enforces that $x$ and $\tilde{x}$ differ only in coordinate $i$. This construction shows that the reverse process also evolves by flipping one coordinate at a time, but the rates are adjusted by the relative likelihood of the target state compared to the current state under the reverse marginal $q_{T-t}$. This ratio defines the discrete-state score function, which is central to training the reverse-time model. From the structure of the rate matrix as defined above, Equation (5) can be compactly written as follows:
@@ 3.4 Problem Formulation
Let the discrete-state score function be approximated by a parameterized family of functions $\mathcal{F}_{\Theta}=\{\hat{s}_{\theta}:\theta\in\Theta\}$, where each $\hat{s}_{\theta}:[S]^{d}\times[0,T]\to\mathbb{R}^{d(S-1)}$ is typically represented by a neural network with depth $D$ and width $W$. Where $D$ indicates the number of layers in the neural network and $W$ indicates the maximum number of neurons in any layer of the neural network. Given i.i.d. samples $\{x_{i}\}_{i=1}^{n}\sim p_{\text{data}}$ from the data distribution, the score network is trained by minimizing the following time-indexed loss:
@@ 4 Assumptions
$\mathcal{L}_{k}(\theta)$ $=\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\left[D_{I}\left(s_{k}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,k}(\mathbf{x}_{t})\right)\right]dt$ (14)
@@
$\widehat{\mathcal{L}}_{k}(\theta)=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}D_{I}\left(s_{t}(\mathbf{x}_{k,t,i})\,\|\,\hat{s}_{\theta,k}(\mathbf{x}_{k,t,i})\right),$
$\quad t\sim\text{Unif}[kh,(k+1)h]$ (15)
where $n_{k}$ is the sample size at each time step. $x_{k,t,i}$ denotes the forward-diffused sample obtained at a random time $t$ within the $k$-th discretization interval, where $i$ indexes the training example from a dataset of size $n_{k}$. Here, $k$ denotes the discretization step corresponding to the interval $[kh,(k+1)h]$, and $t\sim\mathrm{Unif}[kh,(k+1)h]$ is a uniformly sampled continuous time.
@@
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks (Liu et al., 2022). Prior works in continuous-state diffusion models such as Gupta et al. (2024) and Block et al. (2020) implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in Gupta et al. (2024) and the definition of $\hat{f}$ in Theorem 13 of Block et al. (2020)). This assumption, however, introduces a major limitation for practical implementations, where the exact ERM is not attainable. In contrast, the PL condition allows us to derive sample complexity bounds under realistic optimization dynamics, without requiring exact ERM solutions.
@@
The bounded gradient variance assumption is a standard assumption used in SGD convergence results such as Koloskova et al. (2022) and Ajalloeian and Stich (2020). This assumption holds for a broad class of neural networks under mild conditions, such as Lipschitz-continuous activations (e.g., GELU), bounded inputs, and standard initialization schemes (Allen-Zhu et al., 2019; Du et al., 2019). The smoothness condition ensures that the gradients of the loss function do not change abruptly, which is crucial for stable updates during optimization. This stability facilitates controlled optimization dynamics when using first-order methods like SGD or Adam, ensuring that the learning process proceeds in a well-behaved manner.
@@
The data distribution $p_{\text{data}}$ has full support on $\mathcal{X}$, and there exists a constant $B>0$, such that for all $\mathbf{x}\in\mathcal{X}$, $i\in[d]$, and $x^{i}\neq\hat{x}^{i}\in[S]$,